OpenStudy (anonymous):

I need help with the steps to solve equation 3/4=3/8x-3/2

OpenStudy (anonymous):

ok first do you want to work with fraction to solve it or you want to get ride of them ?

OpenStudy (anonymous):

work with fractions please

OpenStudy (anonymous):

it would life much easier if you get ride of the fraction but since you insist ill explain :D

OpenStudy (anonymous):

thks

OpenStudy (anonymous):

tell me what you got :D

hero (hero):

\[\frac{3}{4}=\frac{3x}{8} -\frac{3}{2}\] 1. Add 3/2 5o both sides. Let me know what you get afterwards.

hero (hero):

Add 3/2 to both sides

OpenStudy (anonymous):

6/4

hero (hero):

When I said to add 3/2 to both sides, I meant for @magepker728 to do it.

OpenStudy (anonymous):

lol 2 1/4

hero (hero):

When you added 3/2 to both sides you should have gotten this: \[\frac{3}{4} +\frac{3}{2}=\frac{3x}{8} \]

hero (hero):

Do you understand how I got that @magepker728

OpenStudy (anonymous):

yeh and what is the product

hero (hero):

The product? Where do you see a 'product'?

OpenStudy (anonymous):

looks like this drink starting to kick lol, i mean the sum :D

OpenStudy (anonymous):

so is this helping me or magepker? just asking

hero (hero):

@magepker728, did you add the left side yet?

OpenStudy (anonymous):

9/4=3x/8

hero (hero):

Oh, sorry @vmsharrai, I'm just now realizing that you're the one who posted the problem.

hero (hero):

Now, cross multiply

OpenStudy (anonymous):

12x/72

hero (hero):

You mean `72 = 12x`

OpenStudy (anonymous):

yes

hero (hero):

Okay, so just divide both sides by 12

OpenStudy (anonymous):

6 thks for your help